Please help me solve this problem:

Suppose your mom pays you to make a new dining room table for apartment. She tells you that the length of the table should be 3 feet more than twice its width. She also tells you that the area of the rectangular table top is 2w^2+12 square feet. What are the length and width of the table? Please explain!


thankyou so much!!!!!!!!

Answer by scott8148(6628)

"length of the table should be 3 feet more than twice its width" ___ L = 2W + 3

area = length times width = (2W + 3)(W) = 2W^2 + 3W

"She also tells you that the area of the rectangular table top is 2w^2+12"

2W^2 + 3W = 2W^2 + 12

3W = 12 ___ W = 4

substituting ___ L = 2(4) + 3 = 11